OpenStudy (gabylovesyou):

@SolomonZelman point-slope form for (3,12) and (-3,30)

OpenStudy (sleepyjess):

Do you know what point-slope form is?

OpenStudy (sleepyjess):

\(y-y_1=m(x-x_1)\)

OpenStudy (gabylovesyou):

y-y1=m(x-x1)

OpenStudy (solomonzelman):

yes, where (x1, y1) is the point, and m is the slope.

OpenStudy (solomonzelman):

use the 2 points to find the slope, first.

OpenStudy (sleepyjess):

Can you find the slope between the 2 points?

OpenStudy (gabylovesyou):

-3

OpenStudy (solomonzelman):